# Feedback loop dependent charge density wave imaging by scanning tunneling spectroscopy

### v1

@@ I Introduction
Here, we discuss specific effects of open versus closed feedback loops on the CDW contrast inversion (CCI) observed in STS conductance maps, which has been demonstrated to inform about the width ($\Delta$) and position ($\epsilon$) of the CDW gap relative to the Fermi level ($E_{F}$)  [8]. We consider the $2a_{0}\times 2a_{0}$ CDW developing below 202 Kelvin in 1T-TiSe2. It is locked to the lattice and well described by a single CDW gap, avoiding the complications that may arise in materials presenting multiple CDW gaps [9, 10]. Most importantly for the present study, $\Delta$ opens below $E_{F}$ in 1T-TiSe2, which, as we will show, results in a more complex CDW contrast as a function of energy, enabling a more complete discussion of the different STS acquisition modes.
@@ II Experimental
Single crystals of 1T-TiSe2 were grown via iodine vapour transport and post-annealed for one week at 650 °C to reduce the amount of self-intercalated Ti. They were cleaved in-situ in ultra-high vacuum at room temperature shortly before mounting them on the STM head. All STM and STS data were measured at 4.5 Kelvin using a SPECS Joule-Thomson STM with a base pressure better than $1\times 10^{-10}$ mbar. The bias voltage was applied to the sample. Tips were mechanically cut from a PtIr wire and conditioned in-situ by Argon sputtering and scanning a clean Ag(111) surface. STS measurements were performed using a lock-in technique with $V_{\text{AC}}=3.54$ mV rms at 413.7 Hz.
@@ III Theoretical model
where $E$ is the quasiparticle energy with respect to $E_{F}$, $\mathit{\Gamma}$ accounts for thermal and non-thermal broadening, $\lambda_{\text{CDW}}$ is the CDW wavelength, and $\rho_{0}$ takes care of the fact that the CDW gap does not open for all momenta in the $(x,y)$ plane.
Fig. 2: Topographic STM patterns of a $2\times 2$ CDW at $V_{b}$ below and above the CDW gap. (a) STM topographies simulated using Eq. (5) at $V_{b}=\pm 5\Delta$ for a CDW gap centered at $E_{F}$ ($\epsilon=0$), including the atomic lattice, and the first and second order CDW components whose phases are set to $32\pi/$ 3 and $\pi$, respectively. (b) Experimental STM topographies measured below and above $\Delta$ on 1T-TiSe2 (adapted from Spera et al [8]). (c), (d) Data in (a) and (b), respectively, filtered to their first and second order CDW components. We refer to the pattern above (below) $E_{F}$ as pattern A (pattern B).
@@
where $\mathbf{k}_{i}$ are the independent wave vectors describing the 3Q CDW of 1T-TiSe2, each with a phase $\phi_{i}$ [14] and $\rho_{at}$ is a spatially modulated and energy-independent LDOS reproducing the atomic modulation.
@@ IV Results
The CDW ground state in the classic Peierls mechanism is described as a lattice distortion and a concomitant local charge redistribution such that the reconstructed LDOS below and above the CDW gap are staggered, as illustrated in Fig. 1 for a 1D CDW. Thus, one expects a contrast inversion between STM and STS maps below and above $\Delta$ [15, 16]. Proper visualization of this contrast inversion in 1T-TiSe2 requires to consider not only the first order CDW Fourier components, but also their differences (i.e. second order Fourier components) [8]. Therefore, we filter all STM and STS maps to the first and second order Fourier components for the analysis as shown in Figs. 2c and d. From here on, we refer to the imaging textures above and below the CDW gap as pattern A and pattern B, respectively.
Fig. 3: Experimental STS data filtered to the first and second order CDW components as a function of the imaging mode: (a) CH-CITS at $V_{b}=-300\,\mathrm{mV}$, (b) CH-CITS at $V_{b}=+150\,\mathrm{mV}$, (c) CC-CITS at $V_{b}=-300\,\mathrm{mV}$, (d) CC-CITS at $V_{b}=+150\,\mathrm{mV}$ and (e) CC-CI. The last column depicts graphically the energy ranges where pattern A (color-coded green) and pattern B (color-coded red) are observed for each imaging mode. Yellow depicts bias ranges where the CDW amplitude is too faint to be unambiguously identified, gray indicates the bias range where no data could be acquired, and the blue arrows mark the gap edges as found in Ref. [8]. For all the CITS modes (a-d), the energy of each 2D cut from the full dataset is specified on top of the corresponding image, and the stabilization biases, marked with a red arrow, are outside the gap. For the CC-CI mode (e) the energy cut coincides with the stabilization bias. Images size: $2.5\times 2.5\leavevmode\nobreak\ \mathrm{nm}^{2}$.
@@
The conclusion at this stage is that only data obtained using CH-CITS enable a direct and unambiguous extraction of the CDW gap amplitude and position in energy (Fig. 3(a,b)) from the contrast inversion across the CDW gap as proposed by Spera et al. [8]. Data obtained using the more common constant current (CDW) imaging modes must be examined much more carefully. In that case, CCIs can be misleading since they may be observed at energies unrelated to a CDW gap or may be totally absent.

### v2

@@ I Introduction
Here, we discuss specific effects of open versus closed feedback loops on the CDW contrast inversion (CCI) observed in STS conductance maps, which has been demonstrated to inform about the width ($\Delta$) and position ($\epsilon$) of the CDW gap relative to the Fermi level ($E_{F}$)  [8]. We consider the commensurate $2a\times 2a$ CDW developing below 202 Kelvin in 1T-TiSe2. It is locked to the lattice and well described by a single CDW gap, avoiding the complications that may arise in materials presenting multiple CDW gaps [9, 10]. Most importantly for the present study, $\Delta$ opens below $E_{F}$ in 1T-TiSe2, which, as we will show, results in a more complex CDW contrast as a function of energy, enabling a more complete discussion of the different STS acquisition modes.
@@ II Experimental
Single crystals of 1T-TiSe2 were grown via iodine vapour transport for 1 month at 590 °C to reduce the amount of self-intercalated Ti. They were cleaved in-situ in ultra-high vacuum at room temperature shortly before mounting them on the STM head. All STM and STS data were measured at 4.5 Kelvin using a SPECS Joule-Thomson STM with a base pressure better than $1\times 10^{-10}$ mbar. The bias voltage was applied to the sample. Tips were mechanically cut from a PtIr wire and conditioned in-situ by Argon sputtering and scanning a clean Ag(111) surface. STS measurements were performed using a lock-in technique with $V_{\text{AC}}=3.54$ mV rms at 413.7 Hz.
@@ III Theoretical model
where $E$ is the quasiparticle energy with respect to $E_{F}$, $\mathit{\Gamma}$ accounts for thermal and non-thermal broadening, $\lambda_{\text{CDW}}$ is the CDW wavelength, $\rho_{0}$ takes care of the fact that the CDW gap does not open for all momenta in the $(k_{x},k_{y})$ plane, and $\Delta$ is the width of the CDW gap, corresponding to the energy difference between the coherence peaks and the gap center, which we assume to be the same for all components in agreement with 1T-TiSe2 being a single gap CDW material.
@@
where $\rho_{at}$ is a spatially modulated and energy-independent LDOS reproducing the atomic modulation and $\mathbf{k}_{i}$ are the independent wave vectors describing the 3Q CDW of 1T-TiSe2, with an angle of 120° between each of them each with a phase $\phi_{i}$ [14].
@@
Fig. 2: Topographic STM patterns of a $2a\times 2a$ CDW at $V_{b}$ below and above the CDW gap. (a) STM topographies simulated using Eq. (5) at $V_{b}=\pm 5\Delta$ for a CDW gap centered at $E_{F}$ ($\epsilon=0$), including the atomic lattice, and the first and second order CDW components whose phases are set to $32\pi/$ 3 and $\pi$, respectively. The $\mathbf{k}_{i}$ wave vectors of the first order components are oriented along $\pi/2+(i-1)2\pi/3$ for $i=1,2,3$. (b) Experimental STM topographies measured below and above $\Delta$ on 1T-TiSe2 (adapted from Spera et al [8]). (c), (d) Data in (a) and (b), respectively, filtered to their first and second order CDW components. We refer to the pattern above (below) $E_{F}$ as pattern A (pattern B).
@@ IV Results
The CDW ground state in the classic Peierls mechanism is described as a lattice distortion and a concomitant local charge redistribution such that the reconstructed LDOS below and above the CDW gap are staggered, as illustrated in Fig. 1 for a 1D CDW. Thus, one expects a contrast inversion between STM and STS maps below and above $\Delta$, which has indeed been observed in quasi-1D [15] and 2D [16, 17] systems. Proper visualization of this contrast inversion in 1T-TiSe2 requires to consider not only the first order CDW Fourier components, corresponding to the doubling of the lattice parameters, but also their differences (i.e. second order Fourier components) [8]. Therefore, we filter all STM and STS maps to the first and second order Fourier components for the analysis as shown in Figs. 2c and d. From here on, we refer to the imaging textures above and below the CDW gap as pattern A and pattern B, respectively.
Fig. 3: Experimental STS data filtered to the first and second order CDW components as a function of the imaging mode: (a) CH-CITS at $V_{b}=-300\,\mathrm{mV}$, (b) CH-CITS at $V_{b}=+150\,\mathrm{mV}$, (c) CC-CITS at $V_{b}=-300\,\mathrm{mV}$, (d) CC-CITS at $V_{b}=+150\,\mathrm{mV}$ and (e) CC-CI. The last column depicts graphically the energy ranges where pattern A (color-coded green) and pattern B (color-coded red) are observed for each imaging mode. Yellow depicts bias ranges where the CDW amplitude is too faint to be unambiguously identified, gray indicates the bias range where no data could be acquired, and the blue arrowsat $-90\pm 70\,\mathrm{mV}$ mark the gap edges as found in Ref. [8]. For all the CITS modes (a-d), the energy of each 2D cut from the full dataset is specified on top of the corresponding image, and the stabilization biases, marked with a red arrow, are outside the gap. For the CC-CI mode (e) the energy cut coincides with the stabilization bias. Images size: $2.5\times 2.5\leavevmode\nobreak\ \mathrm{nm}^{2}$.
@@
The conclusion at this stage is that only data obtained using CH-CITS enable a direct and unambiguous extraction of the CDW gap amplitude and position in energy (Fig. 3(a,b)) from the contrast inversion across the CDW gap as proposed by Spera et al. [8]. Data obtained using the more common constant current imaging modes must be examined much more carefully. In that case, CCIs can be misleading since they may be observed at energies unrelated to a CDW gap or may be totally absent.
V Discussion